# APCoTTA: Continual Test-Time Adaptation for Semantic Segmentation of Airborne LiDAR Point Clouds

### v1

@@ Abstract
Airborne laser scanning (ALS) point cloud segmentation is a fundamental task for large-scale 3D scene understanding. In real-world applications, models are typically fixed after training. However, domain shifts caused by changes in the environment, sensor types, or sensor degradation often lead to a decline in model performance. Continuous Test-Time Adaptation (CTTA) offers a solution by adapting a source-pretrained model to evolving, unlabeled target domains during deployment. Despite its potential, research on ALS point clouds remains limited, facing challenges such as the absence of standardized datasets for ALS CTTA tasks and the risk of catastrophic forgetting and error accumulation during prolonged adaptation. To tackle these challenges, we propose APCoTTA, the first CTTA method tailored for ALS point cloud semantic segmentation. We propose a dynamic trainable layer selection module. This module utilizes gradient information to identify low-confidence layers with near-uniform distributions. These layers are selected for training, and the remaining layers are kept frozen. This design helps retain critical source domain knowledge, mitigating catastrophic forgetting. To further reduce error accumulation, we propose an entropy-based consistency loss. Extremely low-confidence samples often cause unstable gradients and model collapse. By losing such samples based on entropy, we apply consistency loss only to the reliable samples, enhancing model stability. In addition, we propose a random parameter interpolation mechanism, which randomly blends parameters from the selected trainable layers with those of the source model. This approach helps balance target adaptation and source knowledge retention, further alleviating forgetting. Finally, we construct two benchmarks, ISPRSC and H3DC, to address the lack of CTTA benchmarks for ALS point cloud segmentation. Experimental results demonstrate that APCoTTA achieves the best performance on two benchmarks, with mIoU improvements of approximately 9% and 14% over direct inference. The new benchmarks and code are available at https://github.com/Gaoyuan2/APCoTTA.
@@ 1 Introduction
The widespread adoption of helicopters and unmanned aerial vehicle (UAV) platforms has made airborne LiDAR (ALS) data acquisition more flexible, efficient, and cost-effective. These platforms can carry lightweight, high-performance LiDAR sensors that continuously emit laser pulses and record return times and intensities during flight, enabling the rapid generation of high-density, high-accuracy 3D point cloud data. This provides reliable spatial information for applications such as building mapping [58], building heights extraction [61], 3D reconstruction [45, 34], and forest parameter inversion [26, 67, 56]. For these applications, semantic segmentation of ALS point clouds is a key step, enabling the automatic segmentation of surface features such as ground, vegetation, buildings, and cars, thereby providing accurate data support for further analysis.
After training on a specific dataset, the model is deployed with fixed parameters and remains applicable only to those scenarios. In real-world production environments, changes in various factors often induce domain shift, leading to a decline in model performance. First, weather conditions significantly affect ALS point clouds: intense sunlight can cause reflection saturation, resulting in data loss in certain areas, while rain can scatter and absorb laser beams, reducing the number of return points and lowering the signal-to-noise ratio. Second, variations in LiDAR sensors, including emission frequency, scanning patterns, and flight altitude, can cause significant differences in point cloud density and noise levels. Moreover, sensors are susceptible to various types of degradation caused by factors such as vibrations, external lighting conditions, and reflective materials, which can further degrade data quality [9]. Furthermore, variations in terrain structures (e.g., urban high-rises, hills, and farmlands), regional differences in architectural styles, road infrastructure, and public space layouts, as well as seasonal changes in surface features such as vegetation coverage and reflectivity, can all degrade model performance.
These issues are typically addressed through two main approaches: (1) acquiring new raw data with annotations and applying unsupervised domain adaptation (UDA) techniques and (2) continuously retraining the model to adapt to each domain shift. However, both approaches are impractical for remote sensing applications. Due to privacy concerns, access to source domain data may be restricted. Moreover, retraining the model depends on high-quality, high-density annotations [11], and the annotation process is time-consuming and requires significant human effort, computational resources, and storage, leading to high training costs [10].
Test-time adaptation (TTA) [51, 31] has gained significant interest. In this setting, source domain data is inaccessible, and the model adapts to the unlabeled target domain data during online testing. Currently, most TTA methods are focused on 2D images and employ techniques such as pseudo-labeling [4, 12] and entropy minimization [51, 31]. This method is primarily designed for static, single target domains. When the target domain undergoes continuous changes, distribution shifts may generate incorrect pseudo-labeling of unlabeled test data, leading to error accumulation [53, 55]. Furthermore, prolonged adaptation to the target distribution may cause the model to forget source domain knowledge, leading to catastrophic forgetting [44].
Continual Test-Time Adaptation (CTTA) enables the model to continuously adapt to a changing target domain without accessing source domain data, effectively addressing these challenges. To the best of our knowledge, most existing research focuses on 2D images [53, 64, 28], while studies on 3D point cloud remain limited [2, 17]. Common techniques include mean teacher [46], augmentation-averaged predictions [27], contrastive loss [8], restoring source parameters [1], and selecting reliable samples [55]. For 3D point clouds, Cao et al. [2] propose a cross-modal CTTA method called CoMAC, designed for autonomous driving images and point clouds. Jiang et al. [17] introduce a more general CTTA approach, PCoTTA, for object-level point clouds, applicable to tasks such as point cloud reconstruction, denoising, and registration. However, CTTA research on ALS point clouds remains largely unexplored, primarily due to two unresolved problems.:
@@
    The absence of a dedicated benchmark dataset for ALS CTTA tasks, which hinders fair evaluation and comparison of methods;
@@
    Prolonged adaptation to target domains often leads to catastrophic forgetting and error accumulation, resulting in overall performance degradation.
In this paper, we are the first to address these problems by proposing APCoTTA, a novel continual test-time adaptation method for semantic segmentation of ALS point clouds. Recognizing that the pretrained model retains essential source domain knowledge, we aim to preserve most of it by selectively freezing network layers based on gradient norms while allowing a subset of layers to remain trainable to mitigate catastrophic forgetting during adaptation. To address error accumulation caused by incorrect pseudo-labels, we propose an entropy-based consistency loss that filters out low-confidence samples, effectively reducing the propagation of incorrect information. Despite these measures, prolonged adaptation may still lead to catastrophic forgetting. To counter this, we randomly select a subset of trainable parameters and blend them with their counterparts from the source model, thereby preventing overfitting to the target domain and maintaining proximity to the source distribution. Additionally, we construct two robustness benchmark datasets, ISPRSC and H3DC, to address the lack of public CTTA datasets for ALS point cloud semantic segmentation. These benchmarks enable systematic evaluation of segmentation performance under CTTA scenarios and provide standardized evaluation protocols for future research.
To summarize, our contributions are fourfold:
@@
    We propose the first continual test-time adaptation (CTTA) method, APCoTTA, for airborne LiDAR point cloud semantic segmentation.
@@
    We introduce the first robustness benchmark for airborne LiDAR point cloud segmentation, including ISPRSC and H3DC, which incorporates seven common corruption types reflecting adverse weather, sensor measurement bias, and complex real-world noise.
@@
    To mitigate catastrophic forgetting and error accumulation during adaptation, we employ dynamic selection of trainable layers, entropy-based consistency loss, and a randomized parameter interpolation mechanism.
@@
    Extensive evaluations on multiple challenging datasets demonstrate that our method, APCoTTA, significantly improves CTTA performance in segmentation tasks, validating its effectiveness and generalizability.
The remainder of this paper is organized as follows: Section 2 surveys prior studies in point cloud unsupervised domain adaptation, (continual) test time adaptation, and corruptions in 3D semantic segmentation. Our proposed framework is described in Section 3. Section 4 presents comparative experiments, ablation studies, and hyperparameter sensitivity analysis. Section 5 concludes the work and discusses potential directions for future research.
@@ 2.1 Point Cloud Unsupervised Domain Adaptation
Unsupervised Domain Adaptation (UDA) is to improve model performance on an unlabeled target domain by leveraging labeled data from a source domain, despite the presence of distribution shifts between the two domains. It includes adversarial learning, feature alignment, self-supervised learning, and pseudo-labeling methods.
Wu et al. [57] generate unlimited synthetic data using simulators such as GTA-V and employ Geodesic Correlation Alignment to align the feature distribution differences between the source and target domains. Yi et al. [62] complete sparse LiDAR point clouds and optimize completion quality with local adversarial learning. Semantic segmentation is then performed on the normalized completed point clouds, with results projected back to the target domain. Xiao et al. [59] design PCT (Point Cloud Translator), a point cloud transformation method that reduces domain shift between synthetic and real point clouds by addressing appearance and sparsity differences separately. Rochan et al. [36] addresses cross-domain generalization in LiDAR semantic segmentation using self-supervised learning. Saltori et al. [38] perform data augmentation at both local and global levels and mitigate domain shift by cross-mixing point clouds between dual-branch symmetric networks. Li et al. [25] reduce domain gaps by applying adaptive masked noise to source point clouds and leveraging adversarial training to simulate real noise distributions. Although these methods have achieved good results, they cannot address cases where source domain data is inaccessible due to privacy or other constraints.
@@ 2.2 Test-Time Adaptation
Test-Time Adaptation (TTA) is a method that optimizes a model online during inference using only unlabeled test data to adapt to target domain shifts without requiring access to source data. Compared to UDA, TTA operates without source data and enables real-time model adjustment, making it suitable for scenarios with privacy constraints or dynamic data distributions. TTA has rapidly advanced in the 2D image domain, including classification [51, 31, 65], segmentation [66, 54], and object detection [19, 50, 49].
For 3D point cloud semantic segmentation, Saltori et al. [39] propose an online unsupervised domain adaptation method that requires no source data. It adaptively selects high-confidence pseudo-labels through self-training and propagates reliable labels using geometric feature propagation, enabling efficient online adaptation to new environments. Shin et al. [43] enable rapid model adaptation across domains (e.g., different sensors, synthetic-to-real, day-to-night) without source data through intra-modal pseudo-label optimization (Intra-PG) and inter-modal pseudo-label refinement (Inter-PR). Wang et al. [52] propose a test-time adaptation (TTA) method for addressing domain shift in geospatial point cloud semantic segmentation. It adapts to new domains by dynamically updating batch normalization statistics and improves classification accuracy through information maximization and reliability-constrained pseudo-labeling. HGL [69] enhances model generalization across different scenarios through geometric learning at three levels: local (point-level), global (object-level), and temporal (frame-level). In addition, there are other studies on point clouds, such as object detection [6, 5], point cloud classification [30, 7, 42], upsampling [14], and registration [13].
@@ 2.3 Continual Test-Time Adaptation
CoTTA [53], the first CTTA method, builds on self-training with a weight-averaged teacher and augmented mean predictions to improve pseudo-label quality and reduce error accumulation. It also restores partial network weights to source parameters to preserve knowledge and prevent forgetting. DSS [55] dynamically filters high- and low-quality pseudo-labels using adaptive thresholds. It employs positive learning to reinforce high-confidence samples and negative learning to mitigate the impact of low-quality pseudo-labels, reducing error accumulation. Zhu et al. [68] dynamically store high-confidence samples for replay training to reduce error accumulation. Additionally, they maintain inter-class topology using a class relationship graph (CRG) to prevent catastrophic forgetting.
However, research on CTTA for point clouds remains limited. Cao et al. [2] propose CoMAC, a method designed for real-world scenarios with continually changing target domains in multimodal semantic segmentation. It employs a two-stage modality-aware pseudo-labeling mechanism to adaptively select the more reliable modality and suppress noise. Additionally, it constructs class-wise momentum queues to dynamically update class centers and incorporates pseudo-source feature replay to mitigate catastrophic forgetting. PCoTTA [17] prevents catastrophic forgetting by balancing the distance between source prototypes and learnable prototypes. It also employs a Gaussian distribution to align target data features with source domain features, reducing error accumulation. Although these studies have explored CTTA for point clouds, research on ALS point cloud semantic segmentation remains lacking.

### v2

@@ Abstract
Airborne laser scanning (ALS) point cloud semantic segmentation is a fundamental task for large-scale 3D scene understanding. Fixed models deployed in real-world scenarios often suffer from performance degradation due to continuous domain shifts caused by environmental and sensor changes. Continuous Test-Time Adaptation (CTTA) enables adaptation to evolving unlabeled domains, but its application to ALS point clouds remains underexplored, hindered by the lack of benchmarks and the risks of catastrophic forgetting and error accumulation. To address these challenges, we propose APCoTTA (ALS Point cloud Continuous Test-Time Adaptation), a novel CTTA framework tailored for ALS point cloud semantic segmentation. APCoTTA consists of three key components. First, we adapt a gradient-driven layer selection mechanism for ALS point clouds, selectively updating low-confidence layers while freezing stable ones to preserve source knowledge and mitigate catastrophic forgetting. Second, an entropy-based consistency loss discards unreliable samples and enforces consistency regularization solely on reliable ones, effectively reducing error accumulation and improving adaptation stability. Third, a random parameter interpolation mechanism stochastically blends adapted parameters with source model parameters, further balancing target adaptation and source knowledge retention. Finally, we construct two benchmarks, ISPRSC and H3DC, to address the lack of CTTA benchmarks for ALS point cloud segmentation. Extensive experiments demonstrate that APCoTTA achieves superior performance on both benchmarks, improving mIoU by approximately 9% and 14% over direct inference. The new benchmarks and code are available at https://github.com/Gaoyuan2/APCoTTA.
@@ 1 Introduction
Advances in helicopter and unmanned aerial vehicle (UAV) platforms have enabled flexible and efficient acquisition of airborne LiDAR (ALS) point clouds. These data provide critical spatial information for applications including mapping [48, 51], reconstruction  [34, 26], and forest parameter estimation [21, 55, 46]. Semantic segmentation of ALS point clouds performs point-wise semantic labeling, classifying points into categories such as terrain, vegetation, and buildings, which supports downstream analysis and applications.
ALS point cloud semantic segmentation models are often deployed in a fixed-parameter manner after training, which restricts their generalization to the source domain and closely related scenes. This deployment paradigm conflicts with the inherent continuous spatiotemporal variability of ALS data in remote sensing. Spatially, as the flight platform moves, ALS point clouds are acquired in a streaming manner, resulting in progressively changing data distributions, for example, when the surveyed area gradually transitions from urban regions to rural landscapes. Temporally, repeated surveys of the same region further introduce persistent domain shifts, particularly in time-series observations across different seasons, where variations in vegetation phenology, illumination, and surface conditions continuously alter point cloud characteristics. Furthermore, weather and sensor-related factors further intensify these continuous shifts. Adverse weather conditions, such as strong sunlight, rain, or fog, can substantially affect return density and increase noise levels, leading to unstable point cloud characteristics. Differences in LiDAR sensor configurations and flight settings, including emission frequency, scanning patterns, and flight altitude, introduce additional variability in point density and geometric accuracy. Over time, gradual sensor degradation caused by mechanical vibrations and hardware aging further alters measurement stability and data quality [6]. Together, these factors create non-stationary and progressively evolving distributions, posing significant challenges to fixed-parameter models in practical semantic segmentation of ALS point clouds.
To address evolving domain distribution shifts, existing studies primarily adopt three strategies: model retraining, Unsupervised Domain Adaptation (UDA), and Test-Time Adaptation (TTA). However, model retraining necessitates high-quality, dense 3D annotated data [8], incurring prohibitive costs [7]. While UDA methods alleviate the demand for target domain annotations, they still necessitate access to source domain data during the adaptation process, which is often impractical in remote sensing applications constrained by data privacy [8]. Although TTA circumvents the dependency on source domain data and expensive annotations, existing methods [41, 25] are typically predicated on the assumption that the “target domain distribution remains static”, relying on self-supervised objectives such as entropy minimization [41, 25] or pseudo-labeling [4, 9] for optimization. In ALS scenarios, however, continuous spatiotemporal shifts, sensor heterogeneity, and variations in observation conditions render this assumption untenable, leading to a sharp decline in the reliability of pseudo-labels and inducing error accumulation [43, 45]. Furthermore, long-term continual adaptation to dynamic target distributions renders the model prone to a gradual loss of source domain knowledge, ultimately triggering the issue of catastrophic forgetting [35].
These limitations highlight the need for Continual Test-Time Adaptation (CTTA) tailored to ALS point clouds, which aims to enable continuous adaptation to evolving target domains without access to source data. To the best of our knowledge, most existing CTTA research focuses on 2D image tasks [43, 53, 23], while studies on 3D data remain limited and are primarily oriented toward autonomous driving point clouds [2] and object-level point cloud applications [12]. Common CTTA techniques include mean-teacher training [36], augmentation-averaged predictions [22], contrastive regularization [5], source parameter restoration [1], and reliable sample selection [45]. Although it may seem intuitive to transfer image-based CTTA methods (e.g., CoTTA [43]) to ALS point cloud segmentation, such direct application is challenging due to fundamental differences in data characteristics and learning mechanisms.
Regarding data characteristics, unlike 2D images where domain shifts typically manifest as illumination or style variations on fixed pixel grids, ALS point clouds are inherently unstructured, and their distributions are continuously affected by scanning patterns, weather conditions, and sensor degradation. These factors directly alter geometric structures and point densities, rather than producing simple appearance-level shifts as in images. Consequently, key assumptions underlying image-based CTTA often break down in point clouds. For example, the common assumption that augmentations should preserve prediction consistency is unreliable for point clouds, since typical augmentations can significantly change local geometry and density, leading to unstable features and inconsistent outputs. In terms of learning mechanisms, image networks operate on regular pixel grids with relatively stable feature statistics, enabling many CTTA methods to rely on techniques such as BatchNorm statistical alignment for robust adaptation. In contrast, point cloud networks process unordered, sparse, and highly density-variant data, where neighborhood structures change dynamically with sampling conditions, making feature statistics far more unstable than those of images. As a result, directly applying image-based CTTA methods to point clouds often leads to unstable adaptation. For instance, full-parameter updates, commonly adopted in image CTTA, are prone to overfitting and catastrophic forgetting in ALS point cloud scenarios. Moreover, research on ALS CTTA is hindered by the lack of dedicated benchmark datasets and evaluation protocols. Unlike image CTTA, no standardized benchmark currently exists for continuous adaptation of large-scale airborne LiDAR point clouds, making systematic study and fair comparison difficult. Consequently, two key challenges remain unexplored for ALS CTTA.
@@
    Existing CTTA methods are primarily designed for 2D imagery and are ill-suited for ALS point clouds, where direct application often leads to unstable adaptation, error accumulation, and catastrophic forgetting.
@@
    The absence of a standardized benchmark for ALS CTTA prevents rigorous evaluation and comparison of methods under continuous domain shifts.
In this paper,we address these challenges within the specific context of ALS point clouds by proposing APCoTTA, a novel continual test-time adaptation framework. To address catastrophic forgetting, we investigate the transferability of image-based parameter-efficient fine-tuning strategies to ALS point clouds and find that full-parameter updates are unsuitable, while gradient norms in point cloud networks reliably reflect geometric stability. Inspired by this, and leveraging the insights from  [11, 23] that the magnitude of gradients relative to a uniform distribution quantifies the model’s familiarity with data, we introduce a Dynamic Trainable Layer Selection (DSTL) module specifically designed for point cloud networks, which freezes stable layers and updates only geometry-sensitive ones to effectively mitigate catastrophic forgetting. To address error accumulation caused by incorrect pseudo-labels, we propose an Entropy-Based Consistency Loss (EBCL). Unlike generic consistency losses, EBCL explicitly filters out incorrect pseudo-labels generated from low-confidence samples in ALS data, ensuring the model learns strictly from reliable samples. Despite these measures, prolonged adaptation may still lead to catastrophic forgetting. To counter this, we propose a Randomized Parameter Interpolation (RPI) mechanism. Instead of the abrupt ‘hard’ resets used in prior works, RPI employs a ‘soft’ parameter fusion strategy to provide a smoother regularization trajectory for continuous terrain adaptation. Additionally, we construct two robust benchmark datasets, ISPRSC and H3DC, based on the ISPRS [29] and H3D [15] datasets, to address the lack of publicly available CTTA datasets for ALS point cloud semantic segmentation. These benchmarks enable systematic evaluation of segmentation performance under CTTA scenarios and provide standardized evaluation protocols for future research. To summarize, our contributions are fourfold:
@@
    We introduce two robustness benchmarks for airborne LiDAR: ISPRSC and H3DC (Sec. 4.1). These benchmarks simulate seven common corruption types, including weather-related noise (e.g., strong sunlight) and sensor-level measurement biases. This fills the gap in standardized datasets for evaluating the robustness of ALS point cloud segmentation under adverse real-world conditions.
@@
    We propose APCoTTA (Fig. 1), a novel CTTA framework specifically tailored for airborne LiDAR point cloud segmentation (Sec. 3). Unlike generic CTTA frameworks, APCoTTA is designed to explicitly address the severe continuous distribution shifts and low-confidence samples in ALS data, ensuring stable and reliable adaptation during deployment without requiring source data.
@@
    We propose three synergistic modules to address the core challenges of ALS adaptation. Specifically, the DSTL module identifies layers sensitive to domain shifts via gradient norms to balance target adaptation with source knowledge retention, effectively mitigating model collapse on sparse data (Sec. 3.2). Complementing this, the EBCL module filters out incorrect predictions caused by low-confidence samples, forcing the model to learn strictly from reliable samples (Sec. 3.3) . Finally, the RPI module mechanism employs a ‘soft’ parameter fusion strategy instead of the abrupt ‘hard’ weight resets used in prior works, providing a smoother regularization trajectory for continuous terrain adaptation (Sec. 3.4).
@@
    Extensive evaluations demonstrate that APCoTTA achieves superior performance (Sec. 4.5). Our method achieves mIoU improvements of approximately 9% and 14% over direct inference on the ISPRSC and H3DC benchmarks, respectively. The results validate the effectiveness and generalizability of our approach in handling diverse and evolving domain shifts in ALS scenarios.
@@ 2.1 Point Cloud Unsupervised Domain Adaptation
Unsupervised Domain Adaptation (UDA) aims to improve model performance on an unlabeled target domain by leveraging labeled data from a source domain, despite the presence of distribution shifts between the two domains. It includes adversarial learning, feature alignment, self-supervised learning, and pseudo-labeling methods.
Wu et al. [47] generate diverse synthetic data using simulators such as GTA-V and employ Geodesic Correlation Alignment to align the feature distribution differences between the source and target domains. Yi et al. [52] complete sparse LiDAR point clouds and optimize completion quality with local adversarial learning. Semantic segmentation is then performed on the normalized completed point clouds, with results projected back to the target domain. Xiao et al. [49] design PCT (Point Cloud Translator), a point cloud transformation method that reduces domain shift between synthetic and real point clouds by addressing appearance and sparsity differences separately. Rochan et al. [28] address cross-domain generalization in LiDAR semantic segmentation using self-supervised learning. Saltori et al. [30] perform data augmentation at both local and global levels and mitigate domain shift by cross-mixing point clouds between dual-branch symmetric networks. Li et al. [19] reduce domain gaps by applying adaptive masked noise to source point clouds and leveraging adversarial training to simulate real noise distributions. Although these methods have achieved promising performance, they face challenges in cases where source domain data is inaccessible due to privacy or other constraints.
@@ 2.2 Test-Time Adaptation
Test-Time Adaptation (TTA) refers to a paradigm that optimizes a model online during inference using only unlabeled test data to adapt to target domain shifts without requiring access to source data. Compared to UDA, TTA operates without source data and enables real-time model adjustment, making it suitable for scenarios with privacy constraints or dynamic data distributions. TTA has rapidly advanced in the 2D image domain, including classification [41, 25, 54], segmentation [56, 44], and object detection [14, 40, 39].
For 3D point cloud semantic segmentation, Saltori et al. [31] propose an online unsupervised domain adaptation method that does not require access to source data. It adaptively selects high-confidence pseudo-labels through self-training and propagates reliable labels using geometric feature propagation, enabling efficient online adaptation to new environments. Shin et al. [33] enable rapid model adaptation across domains (e.g., different sensors, synthetic-to-real, day-to-night) without source data through intra-modal pseudo-label optimization (Intra-PG) and inter-modal pseudo-label refinement (Inter-PR). HGL [58] enhances model generalization across different scenarios through geometric learning at three levels: local (point-level), global (object-level), and temporal (frame-level). Most recently, Wang et al. [42] propose a TTA method tailored for addressing domain shift in geospatial point cloud semantic segmentation. It adapts to new domains by dynamically updating batch normalization statistics and improves classification accuracy through information maximization and reliability-constrained pseudo-labeling. However, this method primarily focuses on the TTA setting involving a single, static target domain, and lacks mechanisms to counter the catastrophic forgetting and error accumulation inherent in continual adaptation scenarios. Unlike Wang et al., our APCoTTA incorporates specifically designed modules to mitigate the issues of catastrophic forgetting and error accumulation that arise during the model’s long-term adaptation to unlabeled target domains.
@@ 2.3 Continual Test-Time Adaptation
CoTTA [43], a pioneering CTTA method, builds on self-training with a weight-averaged teacher and augmented mean predictions to improve pseudo-label quality and reduce error accumulation. It also restores partial network weights to source parameters to preserve knowledge and prevent forgetting. Zhu et al. [57] dynamically store high-confidence samples for replay training to reduce error accumulation. Additionally, they maintain inter-class topology using a class relationship graph (CRG) to prevent catastrophic forgetting. PALM [23] introduces a pseudo-label-free adaptive learning rate strategy that uses the gradient norm of the KL divergence to measure prediction uncertainty and select trainable layers while freezing the rest. It then dynamically adjusts layer-wise learning rates based on parameter sensitivity to support adaptation in continuously changing environments. Inspired by this work, we use gradient norms to identify and freeze stable layers. Unlike PALM, which adjusts layer-wise learning rates, our design is specifically tailored for airborne point clouds and mitigates error accumulation and catastrophic forgetting through entropy and parameter weighting rather than learning rate modulation.
However, research on CTTA for point clouds remains relatively underexplored. Cao et al. [2] propose CoMAC, a method designed for real-world scenarios with continually changing target domains in multimodal semantic segmentation. It employs a two-stage modality-aware pseudo-labeling mechanism to adaptively select the more reliable modality and suppress noise. Additionally, it constructs class-wise momentum queues to dynamically update class centers and incorporates pseudo-source feature replay to mitigate catastrophic forgetting. PCoTTA [12] prevents catastrophic forgetting by balancing the distance between source prototypes and learnable prototypes. It also employs a Gaussian distribution to align target data features with source domain features, reducing error accumulation. Although these studies have explored CTTA for point clouds, research on ALS point cloud semantic segmentation remains lacking.